# Generative Modeling of Neural Dynamics via Latent Stochastic Differential Equations

### v2

@@ Abstract
Understanding how neural circuits implement computation remains a central challenge in neuroscience, requiring models that can unite mechanistic interpretability with the complexity of real neural data. Current approaches face a fundamental trade-off: mechanistic biophysical and phenomenological models are difficult to fit to single-trial data, while flexible machine learning models lack interpretability and theoretical grounding. We propose a probabilistic framework that addresses this challenge by combining domain-specific mathematical models with neural networks through stochastic differential equations (SDEs). In this framework, physiological recordings are viewed as discrete-time partial observations of an underlying continuous-time stochastic dynamical system that implements computations through its state evolution. We employ coupled SDEs with differentiable drift and diffusion functions and use variational inference to infer system states and parameters. This formulation enables seamless integration of existing mathematical models from the literature, neural networks, or hybrid combinations to learn and compare different modeling approaches. We demonstrate this framework by developing a generative model that combines coupled oscillators with neural networks to capture latent population dynamics from single-cell recordings. Evaluation across three neuroscience datasets spanning different species, brain regions, and behavioral tasks shows that these hybrid models achieve competitive performance in predicting stimulus-evoked neural and behavioral responses compared to sophisticated black-box approaches, while requiring an order of magnitude fewer parameters, providing principled uncertainty estimates, and offering interpretable mathematical descriptions of neural dynamics.
@@ Methods
where $\tilde{x}_{0}\sim\mathcal{Q}_{0}$ and $c(t)=\sum_{i=1}^{T}\xi_{\phi}({y}_{i},{b}_{i})\psi_{i}(t)$ is a continuous path encoding the observations which we refer to a context, with $\xi_{\phi}\colon\mathbb{R}^{d_{y}\times d_{b}\times d_{v}}\to\mathbb{R}^{d_{c}}$ being a modality-specific observation encoder. We define the distribution of the initial condition as $\mathcal{Q}_{0}=\mathcal{N}(\alpha_{\phi}({c}_{1:n}),\beta_{\phi}({c}_{1:n}))$, where $\alpha_{\phi}$ and $\beta_{\phi}$ are neural networks, and ${c}_{1:n}=\{{c}_{1},\ldots,{c}_{n}\}$ are discrete samples of the continuous-time context $c(t)$. Here, $\nu_{\phi}\colon\mathbb{R}^{d_{x}}\times\mathbb{R}^{d_{u}}\times\mathbb{R}^{d_{c}}\to\mathbb{R}^{d_{x}}$ is a neural network representing the drift term of the augmented SDE. Note that while the drift term $\nu_{\phi}$ is different from that of the generative model, the diffusion term $\sigma_{\theta}$ is shared with the generative SDE.
@@
$\log p(\mathcal{Y},\mathcal{B}\mid\mathcal{V})\geq\mathbb{E}_{\tilde{x}}\left[\sum_{i=1}^{T}\log p({y}_{i}|\tilde{x}_{i})+\sum_{i=1}^{T}\log p({b}_{i}|\tilde{x}_{i})\right]-D_{KL}(\mathcal{Q}_{0}||\mathcal{P}_{0})-D_{KL}(\mathcal{Q}_{\tau}||\mathcal{P}_{\tau})$ (6)
@@
$b_{i}\sim\mathcal{N}(\rho_{\theta}^{n_{1}}(x_{i}),\rho_{\theta}^{n_{2}}(x_{i}))$ (11)
where $\rho_{\theta}^{n_{1}}:\mathbb{R}^{d_{x}}\to\mathbb{R}^{d_{b}}$ and $\rho_{\theta}^{n_{2}}:\mathbb{R}^{d_{x}}\to\mathbb{R}^{d_{b}}$ are neural networks that map the latent state to the mean and variance parameters of the observation distribution, respectively.
@@
$b_{i}\sim\text{Categorical}(\rho^{b}_{\theta}(x_{i}))$ (12)
where $\rho^{b}_{\theta}:\mathbb{R}^{d_{x}}\to\Delta^{d_{b}}$ is a neural network followed by a softmax activation that maps the latent state to a probability vector over $d_{b}$ discrete categories.
@@ Results
Fig 3: Generative modeling of neural and behavioral data via latent coupled oscillators across three different datasets. (a-c) Perturbed reach task: The generative model takes as input the target location and manipulandum forces/torques and is tasked with predicting neural responses of 64 Neurons in Area 2 and the behavioral response of the monkey as measured via the cursor position. The inferred frequency of the coupled oscillators as well as the coupling strength (for a sample trial) for one training run is show in c. (d-f) Visual decision-making task: Input consists of contrast levels as well as the timing of presentation and the go cue. The model is trained to predict both neural activity across multiple brain regions and wheel velocity. (g,h,k) Delayed reach task: Unlike the other tasks, this model receives only preparatory neural activity before movement onset and is trained to predict subsequent neural activity in the dorsal pre-motor (PMd) and primary motor cortex (M1) during movement execution. Abbreviations: MOp: primary motor area, LSc: lateral sensory cortex, PT: posterior thalamus, CP: caudoputamen, LSr: Lateral sensory rostral area, PMd: dorsal pre-motor cortex, M1: primary motor cortex. Monkey and mouse illustrations adapted from scidraw.io ($C$-BY 4.0)
@@
Beyond quantitative performance, oscillator-based models offer interpretable parameterizations that connect to established neuroscience concepts. We visualize the inferred natural frequencies and coupling strengths learned by the model in Fig. 3. In both motor tasks —the perturbed reach task with recordings from somatosensory cortex and the delayed reach task with recordings from two motor cortical areas- we observe that coupling strength increases over time during movement preparation and execution (Fig.3c,k). This pattern is consistent with theories of neural communication through enhanced synchronization [26], where increasing oscillatory coupling may facilitate information integration across neural populations during coordinated motor output. In contrast, the visual decision-making task shows a markedly different pattern (Fig.3f). Here, coupling strength decreases over time, with two distinct drops coinciding with the stimulus presentation and go cue. This temporal modulation may reflect dynamic routing of information[26], where coupling is selectively reduced to gate irrelevant signals or reallocate processing resources as task demands shift from sensory encoding to decision formation and action selection. We also observe systematic differences in natural frequencies across brain regions. Notably, oscillators corresponding to M1 exhibit higher average frequencies than those in PMd during the delayed reach task (Fig.3h). This frequency gradient is consistent with the hierarchical organization of motor cortex, where M1 operates at faster timescales than PMd, which is more involved in movement planning [73, 13]. Such frequency differences may reflect distinct computational roles: higher frequencies in M1 may support rapid sensorimotor transformations required for real-time motor control, while lower frequencies in PMd may be suited for integrating information over longer timescales during motor preparation. Important note on interpretation: While these learned parameters show relative consistency across model initializations and align with neuroscience literature, we emphasize that our method does not provide theoretical guarantees for parameter identifiability—a well-known challenge in nonlinear dynamical systems [54]. Therefore, we present these findings as hypothesis-generating observations that suggest our model captures functionally meaningful structure, rather than as definitive claims about the underlying neural mechanisms. The value of these parameterizations lies in providing neuroscience-motivated, interpretable descriptions of population dynamics that can guide future experimental investigations.
@@ 2 Implementation details
Learning rate 1e-3
Max n_epochs 2000

### v3

@@ Abstract
Understanding how neural circuits implement computation remains a central challenge in neuroscience, requiring models that can unite mechanistic interpretability with the complexity of real neural data. Current approaches face a fundamental trade-off: mechanistic biophysical and phenomenological models are difficult to fit to single-trial data, while flexible machine learning models lack interpretability and theoretical grounding. We propose a probabilistic framework that addresses this challenge by combining domain-specific mathematical models with neural networks through stochastic differential equations (SDEs). In this framework, physiological recordings are viewed as discrete-time partial observations of an underlying continuous-time stochastic dynamical system that implements computations through its state evolution. We employ coupled SDEs with differentiable drift and diffusion functions and use variational inference to infer system states and parameters. This formulation enables seamless integration of existing mathematical models from the literature, neural networks, or hybrid combinations to learn and compare different modeling approaches. We demonstrate this framework by developing a generative model that combines coupled oscillators with neural networks to capture latent population dynamics from single-cell recordings. Evaluation across three neuroscience datasets spanning different species, brain regions, and behavioral tasks shows that these hybrid models achieve competitive performance in predicting stimulus-evoked neural and behavioral responses compared to sophisticated black-box approaches, while requiring an order of magnitude fewer parameters, providing principled uncertainty estimates, and offering interpretable mathematical descriptions of neural dynamics. [^1] The code is publicly available at https://github.com/elgazzarr/LdX
@@ Methods
where $\tilde{x}_{0}\sim\mathcal{Q}_{0}$ and $c(t)=\sum_{i=1}^{T}\xi_{\phi}({y}_{i},{b}_{i},{v}_{i})\psi_{i}(t)$ is a continuous path encoding the observations which we refer to as context, with $\xi_{\phi}\colon\mathbb{R}^{d_{y}\times d_{b}\times d_{v}}\to\mathbb{R}^{d_{c}}$ being a modality-specific observation encoder. We define the distribution of the initial condition as $\mathcal{Q}_{0}=\mathcal{N}(\alpha_{\phi}({c}_{1:n}),\beta_{\phi}({c}_{1:n}))$, where $\alpha_{\phi}$ and $\beta_{\phi}$ are neural networks, and ${c}_{1:n}=\{{c}_{1},\ldots,{c}_{n}\}$ are discrete samples of the continuous-time context $c(t)$. Here, $\nu_{\phi}\colon\mathbb{R}^{d_{x}}\times\mathbb{R}^{d_{u}}\times\mathbb{R}^{d_{c}}\to\mathbb{R}^{d_{x}}$ is a neural network representing the drift term of the augmented SDE. Note that while the drift term $\nu_{\phi}$ is different from that of the generative model, the diffusion term $\sigma_{\theta}$ is shared with the generative SDE.
@@
$\log p(\mathcal{Y},\mathcal{B}\mid\mathcal{V})\geq\mathbb{E}_{\tilde{x}}\left[\sum_{i=1}^{T}\log p({y}_{i}|\tilde{x}_{i})+\sum_{i=1}^{T}\log p({b}_{i}|\tilde{x}_{i})\right]-D_{KL}(\mathcal{Q}_{0}\|\mathcal{P}_{0})-D_{KL}(\mathcal{Q}_{\tau}\|\mathcal{P}_{\tau})$ (6)
@@
$b_{i}\sim\mathcal{N}(\rho_{\theta}(x_{i}),\rho_{\theta}(x_{i}))$ (11)
where $\rho_{\theta}\colon\mathbb{R}^{d_{x}}\to\mathbb{R}^{d_{b}}$ is a neural network that maps the latent state to the mean and variance parameters of the observation distribution, respectively.
@@
$b_{i}\sim\text{Categorical}(\rho_{\theta}(x_{i}))$ (12)
@@ Results
Fig 3: Generative modeling of neural and behavioral data via latent coupled oscillators across three different datasets. (a-c) Perturbed reach task: The generative model takes as input the target location and manipulandum forces/torques and is tasked with predicting neural responses of 64 Neurons in Area 2 and the behavioral response of the monkey as measured via the cursor position. The inferred frequency of the coupled oscillators as well as the coupling strength (for a sample trial) for one training run is show in c. (d-f) Visual decision-making task: Input consists of contrast levels as well as the timing of presentation and the go cue. The model is trained to predict both neural activity across multiple brain regions and wheel velocity. (g,h,k) Delayed reach task: Unlike the other tasks, this model receives only preparatory neural activity before movement onset and is trained to predict subsequent neural activity in the dorsal pre-motor (PMd) and primary motor cortex (M1) during movement execution. Abbreviations: MOp: primary motor area, LSc: lateral sensory cortex, PT: posterior thalamus, CP: caudoputamen, LSr: Lateral sensory rostral area, PMd: dorsal pre-motor cortex, M1: primary motor cortex. Monkey and mouse illustrations adapted from scidraw.io ($C$-BY 4.0).
@@
Beyond quantitative performance, oscillator-based models offer interpretable parameterizations that connect to established neuroscience concepts. We visualize the inferred natural frequencies and coupling strengths learned by the model in Fig. 3. In both motor tasks – the perturbed reach task with recordings from somatosensory cortex and the delayed reach task with recordings from two motor cortical areas – we observe that coupling strength increases over time during movement preparation and execution (Fig.3c,k). This pattern is consistent with theories of neural communication through enhanced synchronization [26], where increasing oscillatory coupling may facilitate information integration across neural populations during coordinated motor output. In contrast, the visual decision-making task shows a markedly different pattern (Fig.3f). Here, coupling strength decreases over time, with two distinct drops coinciding with the stimulus presentation and go cue. This temporal modulation may reflect dynamic routing of information[26], where coupling is selectively reduced to gate irrelevant signals or reallocate processing resources as task demands shift from sensory encoding to decision formation and action selection. We also observe systematic differences in natural frequencies across brain regions. Notably, oscillators corresponding to M1 exhibit higher average frequencies than those in PMd during the delayed reach task (Fig.3h). This frequency gradient is consistent with the hierarchical organization of motor cortex, where M1 operates at faster timescales than PMd, which is more involved in movement planning [73, 13]. Such frequency differences may reflect distinct computational roles: higher frequencies in M1 may support rapid sensorimotor transformations required for real-time motor control, while lower frequencies in PMd may be suited for integrating information over longer timescales during motor preparation. Important note on interpretation: While these learned parameters show relative consistency across model initializations and align with neuroscience literature, we emphasize that our method does not provide theoretical guarantees for parameter identifiability—a well-known challenge in nonlinear dynamical systems [54]. Therefore, we present these findings as hypothesis-generating observations that suggest our model captures functionally meaningful structure, rather than as definitive claims about the underlying neural mechanisms. The value of these parameterizations lies in providing neuroscience-motivated, interpretable descriptions of population dynamics that can guide future experimental investigations.
@@ 2 Implementation details
Learning rate $10^{-3}$
Max epochs 2000